Problem 3: Chlorine gas is made by reacting HCl and O2 to form Cla and H2O. 100.0 mol/h of HCl are fed to the reactor. Air (Nitrogen and Oxygen) is also fed to the process to provide 25.0% XS oxygen. The fractional conversion of HC1 is 0.700. HCl 02 Cl2 H20 a) Draw and completely label a flow chart (since this is a reactive process, label using individual component flow rates). Do a degree of freedom analysis. b) Determine the molar flow rate of air (O2 and N2) into the reactor. c) Calculate the extent of reaction (5 squiggle) using the balanced equation & information given above. d) Determine the molar flow rates of all components leaving the reactor using the extent of reaction e) Determine the molar flow rates of all components leaving the reactor using atomic balances
Answer #1

Degree of freedom = Number of variables- number of independent equations connecting the variables

Known :Flow rate of HCl, unknown : flow rate of air, independent equation: 2 moles of HCl requires 1 mole of Oxygen. This gives the flow rate of air required. 25% excess air is supplied. Hence air supplied can be calculated. Flow rate of products Cl2 and H2O can also be calculated based on the stoichiometry of the reaction.

The reaction is 2HCl+ 1/2O2--àCl2 + H2O

Basis : 100 mol/hr of HCl. As per the reaction, 2 moles of HCl requires 0.5 moles of oxygen.

100 moles/hr requires 100*0.5/2 moles/hr of Oxygen = 25 moles/hr of oxygen.

Air is supplied 25% in excess and air contains 21%O2 and 79% N2. Moles/ hr of air required = 25/0.21 =119 moles/hr of air. Air supplied is= 119*1.25=149 moles/hr. N2 supplied = 149*0.79= 118 moles/hr. Rest is oxygen which is =149-118= 31 moles/hr

HCl converted =0.7, HCl reacted = 100*0.7= 70 moles/hr

Oxygen consumed = 70*0.5/2 = 17.5 moles/hr. Moles of Cl2 formed = 35 moles/hr, H2O formed = 35 moles/hr

Extent of reaction =100* (100-30)/100=70%

Products contains ( moles/hr) : HCl= 30 moles/hr, O2 = 31-17.5= 13.5, N2= 118, Cl2= 35 and H2O= 35

HCl 100 moles/hr Reactor Products contains N2, H20, CI2, HCl and 02 Air containing 118 moles/hr N2 and 31 moles/hr 02